# An ultraslow optical centrifuge with arbitrarily low rotational acceleration

### v1

@@ VI Summary
In summary, we have developed an alternative implementation of the optical centrifuge, which enabled us to lower its rotational acceleration by three orders of magnitude, as compared to the conventional centrifuge design. We have demonstrated the ability to tune the acceleration in the range of $\sim$100\text{\,}\mathrm{M}\mathrm{H}\mathrm{z}\mathrm{/}\mathrm{p}\mathrm{s}$\$, and applied this “ultraslow optical centrifuge” to css molecules in a molecular jet, confirming their successful capture and spinning by the centrifuge field. Using usCFG to investigate the effect of the helium environment on the rotational dynamics of molecules embedded in helium nanodroplets is the subject of current study. In that regard, the ultraslow centrifuge opens exciting new avenues to probe the interaction of rotationally excited molecules with quantum many-body systems.

### v2

@@ VI Summary
In summary, we have developed an alternative implementation of the optical centrifuge, which enabled us to lower its rotational acceleration by three orders of magnitude, as compared to the conventional centrifuge design. We have demonstrated the ability to tune the acceleration in the range of around $100\text{\,}\mathrm{M}\mathrm{H}\mathrm{z}\mathrm{/}\mathrm{p}\mathrm{s}$, and applied this “ultraslow optical centrifuge” to css molecules in a molecular jet, confirming their successful capture and spinning by the centrifuge field. Using usCFG to investigate the effect of the helium environment on the rotational dynamics of molecules embedded in helium nanodroplets is the subject of current study. In that regard, the ultraslow centrifuge opens exciting new avenues to probe the interaction of rotationally excited molecules with quantum many-body systems.